# Axions on a Hyperbolic Ride: Geometric Suppression of CMB Isocurvature and a Blue-Tilted Spectrum

### v1

@@ Abstract
CMB limits on cold-dark-matter isocurvature are often interpreted as excluding the simultaneous realization of high-scale inflation and large QCD axion decay constants in pre-inflationary Peccei–Quinn (PQ) scenarios. We show that this conclusion can be evaded by exploiting field-space geometry. For a minimal complex PQ scalar with a $U(1)$-symmetric potential and nonlinear sigma-model kinetic term $d\sigma^{2}=dR^{2}+f^{2}(R)\,d\theta^{2}$, a curved target-space metric endows the axion fluctuation with a time-dependent geometric mass during inflation, suppressing isocurvature without explicit PQ breaking and without extreme radial displacements. Specializing to a hyperbolic metric $f(R)=L\sinh(R/L)$ with curvature scale $L$, we find that for $R\gtrsim L$ the canonically normalized angular mode can be generically $\mathcal{O}(H_{\rm inf})$-heavy during radial slow-roll, dynamically damping CMB-scale fluctuations while producing a characteristic blue-tilted isocurvature spectrum. As a result, inflationary Hubble scales as large as $H_{\rm inf}\sim 10^{13}\,\mathrm{GeV}$ can be compatible with $f_{a}\sim 10^{14}$$10^{16}\,\mathrm{GeV}$, reopening parameter space usually regarded as excluded. We present numerical benchmarks and a semi-analytic template that relates the scale-dependence of isocurvature to the geometric lever arm $R/L$, providing a direct phenomenological probe on PQ field-space geometry.
@@ I Introduction
In this work we present a minimal, $U(1)$ conserving alternative based on field-space geometry. Curved internal field-space metrics are well known to generate effective masses and characteristic scale-dependence for entropic/isocurvature modes in covariant sigma-model systems [8, 46, 14, 35, 24]. In this Letter we apply this mechanism directly to the PQ field and show that a hyperbolic field-space metric generically endows the angular fluctuation with a geometric, time-dependent effective mass during inflation, without explicit PQ breaking in the axion potential. This dynamically suppresses CMB-scale axion isocurvature, while the evolving mass produces a distinctive blue-tilted isocurvature spectrum on smaller scales. As a result, viable parameter space with $H_{\rm inf}\sim 10^{13}\,{\rm GeV}$ and large $f_{a}$ (e.g. $\lesssim 10^{16}\,{\rm GeV}$) can be reopened without requiring the extreme displacements $R_{\rm inf}/f_{a}\sim 10^{4}$$10^{5}$ or tiny quartic coefficients $\lesssim\mathcal{O}(10^{-16})$. We further provide a compact phenomenological template relating the isocurvature spectral index to the field-space curvature scale, thereby promoting measurements of isocurvature scale-dependence to a direct probe of PQ field-space geometry and enabling targeted searches in future analyses.
The paper is organized as follows. Section II reviews the axion isocurvature constraint, Sec. III introduces the hyperbolic PQ field-space framework, and Sec. IV presents benchmarks and resulting spectra. We conclude in Sec. V. Appendices A, B and C give a UV-motivated origin for the hyperbolic metric, a semi-analytic data-facing template, and relevant derivations respectively.
@@ II The standard QCD axion isocurvature problem
$T_{\rm inst.RH}$ is the instantaneous reheating temperature and the orange solid line marks the upper boundary of the region for PQ restoration post-inflation. In flat-metric quartic PQ models, realizing simultaneously $H_{\rm inf}\simeq 10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{16}\,{\rm GeV}$ typically forces $\kappa\gtrsim 10^{4}$, often into a regime where the spectator/EFT description is difficult to control. This motivates alternative suppression mechanisms.
@@ III Hyperbolic PQ field-space and geometric suppression
Details and derivations in this section are deferred to the App. C. We work in the vanishing-charge background $\dot{\theta}_{0}=0$, for which the homogeneous evolution is purely radial,
@@
Defining the canonically normalized angular fluctuation
$\delta\psi\equiv f(R_{0})\,\delta\theta,$ (13)
one finds
@@
The leading geometric contribution is enhanced by the curvature lever arm $\xi=R_{0}/L$, while the subleading negative term encodes the reduction of transverse masses from negative curvature when $\dot{R}_{0}\neq 0$. In typical slow-roll trajectories the angular mode is initially $\mathcal{O}(H)$-heavy (damping CMB-scale fluctuations) and becomes lighter as $R_{0}$ relaxes and $\dot{R}_{0}$ redshifts. This time dependence generically produces blue-tilted isocurvature: modes exiting earlier (larger scales) are most suppressed, while later-exiting modes (smaller scales) are less suppressed and approach a quasi-plateau once the angular mode becomes effectively light.
@@ IV Benchmark models and isocurvature spectrum
$\Delta_{\theta}^{2}(k)\equiv\frac{k^{3}}{2\pi^{2}}\,|\delta\theta_{k}|^{2},\qquad\delta\theta_{k}=\left.\frac{\delta\psi_{k}}{f(R_{0})}\right|_{\rm late}.$ (21)
@@
We consider three benchmark points B1-B3 (Tab. 1), chosen such that the initial curvature lever arm $\xi_{i}\equiv R_{i}/L$ is $\sim\mathcal{O}(10)$ and CMB-scale isocurvature is suppressed at high $H_{\rm inf}\sim 10^{13}~\rm{GeV}$. The near-degeneracy of B1 and B3 is intentional: in $H_{\rm inf}$ units the spectator PQ system has an approximate rescaling symmetry,
@@ IV.1 Results
Fig. 2 shows the mass history for B1. The radial field slow-rolls toward its vacuum, while the hyperbolic metric induces a geometric axial mass. Initially $m_{\psi}^{2}\sim\mathcal{O}(H_{\rm inf}^{2})$ while $m^{2}_{R}\ll H_{\rm inf}^{2}$, so modes exiting early are damped. As $R_{0}$ decreases and $\dot{R}_{0}$ redshifts, $m_{\psi}^{2}/H^{2}$ falls and the damping weakens for later-exiting modes. In the regime $\delta=\xi(m_{R}/H_{\rm inf})^{2}\lesssim 1$, the dominant behavior is $m_{\psi}^{2}\simeq\xi\,m_{R}^{2}$, i.e. the angular mode is enhanced relative to the flat case by the curvature lever arm $\xi=R/L$, while for $\xi m_{R}^{2}\gtrsim 1$ the negative curvature term $\propto-\dot{R}^{2}$ reduces the effective axial mass.
@@
Fig. 3 shows the resulting conserved superhorizon $\Delta_{\mathcal{S}}^{2}(k)$. We take $k/(a_{i}H_{\rm inf})=10$ to correspond to the CMB pivot scale $k_{\ast}\simeq 0.002\,{\rm Mpc}^{-1}$. Power is strongly suppressed on the largest CMB scales because those modes exit while $m_{\psi}^{2}\gtrsim H_{\rm inf}^{2}$. As $m_{\psi}^{2}/H_{\rm inf}^{2}$ decreases during the subsequent radial evolution, smaller scales are progressively less suppressed. The outcome is a blue-tilted isocurvature spectrum with running over the range of modes that probe the time variation of $m_{\psi}^{2}$, approaching a quasi-plateau once the angular mode becomes effectively light. Equivalently, the same scale-dependence can be understood kinematically from the conversion $\delta\theta=\delta\psi/f(R)$. For $R\gg L$, $f(R)=L\sinh(R/L)$ is exponentially sensitive to $R$, so slow-roll changes in $R$ translate into strong, $k$-dependent suppression for modes exiting while $R$ is evolving.
@@
A key result is that hyperbolic geometry can reopen parameter space that is excluded in the flat-metric case. In our benchmarks, we focus on $H_{\rm inf}=10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{14}\text{--}10^{16}\,{\rm GeV}$, since this lies near current upper limits and is therefore the most demanding regime for isocurvature. In contrast, for $f(R)=R$ and comparable spectator trajectories, the angular mode remains effectively light, and satisfying the CMB bound at $H_{\rm inf}=10^{13}\,{\rm GeV}$ typically requires $R_{i}/f_{a}\sim 10^{4}\text{--}10^{5}$, often implying super-Planckian displacements.
@@
In the slow-roll geometric domain, the hyperbolic metric ties the axial mass to the radial background through $\xi\equiv R/L$, implying a characteristic non-power-law running. For quartic PQ slow-roll, we give the following scale-dependent mass template (see App. B)
@@
where $k_{m}$ is a matching scale at the onset of the geometric regime and $\xi_{m}\equiv R_{m}/L$ controls the shape. A convenient prescription is to fix $k_{m}$ by $n_{\rm iso}(k_{m})=2$, equivalently $m_{\psi}^{2}(k_{m})/H_{\rm inf}^{2}=5/4$, so that for $k\gtrsim k_{m}$ the entire running is governed by the single curvature parameter $\xi_{m}$, which labels the family of trajectories.
The salient point is the functional form. Combining the hyperbolic relation $m_{\psi}^{2}\propto\xi\,m_{R}^{2}\propto R\,m_{R}^{2}$ with a quartic slow-roll trajectory yields
@@ V Conclusion
We have shown that the intrinsic geometry of the PQ field manifold can provide a minimal and symmetry-preserving solution to the standard axion isocurvature problem. For a hyperbolic PQ target-space metric, the angular fluctuation acquires a geometric, time-dependent effective mass during inflation, even in the absence of explicit PQ breaking in the axion potential. This dynamically suppresses CMB-scale isocurvature while generically yielding a new mechanism to generate blue-tilted spectrum. In benchmark realizations with $H_{\rm inf}=10^{13}\,{\rm GeV}$, the mechanism reopens parameter space at large $f_{a}$ that is typically excluded in flat field-space.
Beyond enlarging the viable QCD-axion window for high-scale inflation, our results highlight a broader point: measurements of isocurvature scale-dependence can directly probe field-space curvature. We provided a compact phenomenological template that ties the running of the isocurvature tilt to a single curvature-controlled parameter, enabling model-agnostic tests of the mechanism and straightforward incorporation into future analyses. Several broad directions are motivated by this perspective. On the theory side, it will be important to develop controlled UV realizations of hyperbolic PQ metrics, and extend the framework to more general target spaces (non-constant curvature and/or non-diagonal metrics). On the dynamics side, nontrivial background motion will be particularly interesting (such as $\dot{\theta}\neq 0$ or post-inflationary evolution). Hyperbolic geometry can naturally support large angular velocities, making contact with rotating-field mechanisms such as Affleck-Dine dynamics in baryogenesis ([2]), axiogenesis [22] and kinetic-misalignment in [21], and more generally with additional signatures beyond the power spectrum [51, 28]. Finally, the generic enhancement of power on sub-CMB scales motivates systematic confrontation with small-scale probes, including Lyman-$\alpha$ forest [9], UV luminosity functions of high-$z$ galaxies [55], 21-cm measurements [45], and strong-lensing substructure [23], thereby enabling targeted searches for scale-dependent isocurvature as a window onto PQ field-space geometry.
@@ Appendix A An example of hyperbolic PQ field-space
In our setup the PQ sector is described by a single complex scalar $\Phi$, but we do not assume a flat field-space. At the level of a low-energy effective theory, the most general two-derivative kinetic term takes the form of a nonlinear sigma model,
@@
where $\phi^{a}$ are the two real components of $\Phi$. The PQ symmetry is realized as a $U(1)$ isometry of the target space (shifts of the angular coordinate), but the presence of this isometry does not fix the curvature. We now give a concrete example showing how a hyperbolic geometry can arise.
@@ Appendix B Semi-analytic template for the tilt and its running
Appendix C Theoretical setup
We start from a canonically normalized complex scalar field $\Phi$ with a $U(1)$-symmetric potential $V(|\Phi|)$ and no explicit $U(1)$-breaking terms,
$S=\int d^{4}x\,\sqrt{-g}\left[-g^{\mu\nu}\partial_{\mu}\Phi^{*}\partial_{\nu}\Phi-V\big(|\Phi|\big)\right],$ (45)
where $g_{\mu\nu}$ is the spacetime metric and $g\equiv\det g_{\mu\nu}$. It is convenient to parameterize the complex field in polar coordinates,
$\Phi(x)=\frac{R(x)}{\sqrt{2}}\,e^{i\theta(x)},$ (46)
so that $|\Phi|=R/\sqrt{2}$ and the potential depends only on the radial field, $V=V(R)$.
In terms of $(R,\theta)$, the kinetic term becomes
$g^{\mu\nu}\partial_{\mu}\Phi^{*}\partial_{\nu}\Phi=-\frac{1}{2}g^{\mu\nu}\left(\partial_{\mu}R\,\partial_{\nu}R+R^{2}\,\partial_{\mu}\theta\,\partial_{\nu}\theta\right),$ (47)
and the action can be written in sigma-model form,
$S=\int d^{4}x\,\sqrt{-g}\left[-\frac{1}{2}G_{IJ}(\phi)\,g^{\mu\nu}\partial_{\mu}\phi^{I}\partial_{\nu}\phi^{J}-V(R)\right],$ (48)
where $\phi^{I}=(R,\theta)$ and the internal metric defines the field-space line element {align} dσ^2 =G_IJ(ϕ) dϕ^Idϕ^J
=G_RR dR^2+G_θθ dθ^2 +(G_Rθ+G_θR)dR dθ. For the canonical complex scalar, the nonvanishing components are
$G_{RR}=1,\qquad G_{\theta\theta}=R^{2},\qquad G_{R\theta}=G_{\theta R}=0.$ (49)
C.1 Nontrivial internal field-space metric
[…]
$d\sigma^{2}=G_{IJ}(\phi)\,d\phi^{I}d\phi^{J}=dR^{2}+f^{2}(R)\,d\theta^{2},$ (50)
where $f(R)$ is an arbitrary function of the radial field. For simplicity we restrict to diagonal metrics ($G_{R\theta}=G_{\theta R}=0$). More general non-diagonal metrics can introduce kinetic mixing between $R$ and $\theta$ and lead to richer phenomenology. Absence of ghost instabilities requires a positive-definite metric, in particular $f^{2}(R)>0$. The action becomes {align} S=∫d^4x -g[ -12g^μν(∂_μR ∂_νR +f^2(R) ∂_μθ ∂_νθ)
-V(R)], which reduces to the canonical flat-metric case when $f(R)=R$.
C.2 Background equations of motion
We consider homogeneous background fields,
$R=R(t),\qquad\theta=\theta(t),$ (51)
on an FRW background with $H\equiv\dot{a}/a$. The reduced action is
$S_{\rm bg}=\int dt\,a^{3}(t)\left[\frac{1}{2}\dot{R}^{2}+\frac{1}{2}f^{2}(R)\dot{\theta}^{2}-V(R)\right].$ (52)
The background equations can be written covariantly as
$D_{t}\dot{\phi}^{I}+3H\dot{\phi}^{I}+G^{IJ}V_{,J}=0,$ (53)
[…]
$\Gamma_{\theta\theta}^{R}=-f(R)f^{\prime}(R),\qquad\Gamma_{R\theta}^{\theta}=\Gamma_{\theta R}^{\theta}=\frac{f^{\prime}(R)}{f(R)},$ (54)
with $f^{\prime}(R)=df/dR$. The $R$ equation becomes
$\ddot{R}+3H\dot{R}-f(R)f^{\prime}(R)\,\dot{\theta}^{2}+V_{,R}(R)=0,$ (55)
and the $\theta$ equation is
$\ddot{\theta}+2\frac{f^{\prime}(R)}{f(R)}\dot{R}\,\dot{\theta}+3H\dot{\theta}=0.$ (56)
[…]
$a^{3}(t)\,f^{2}(R)\dot{\theta}=\text{const}.$ (57)
In this work we focus on the vanishing-charge sector,
$\dot{\theta}_{0}=0,\qquad\theta_{0}=\text{const},$ (58)
for which the angular equation is automatically satisfied and the background evolution is purely radial,
$\ddot{R}_{0}+3H\dot{R}_{0}+V_{,R}(R_{0})=0.$ (59)
Thus, in the zero-charge sector the homogeneous background does not depend explicitly on the choice of $f(R)$.
C.3 Fluctuations: radial $\delta R$ and canonical axial $\delta\psi$
We now consider linear fluctuations about the homogeneous background,
$R(t,\bm{x})=R_{0}(t)+\delta R(t,\bm{x}),\qquad\theta(t,\bm{x})=\theta_{0}+\delta\theta(t,\bm{x}),$ (60)
[…]
+12f^2(R_0)(˙δθ^2 -(∇δθ)2a2)], where derivatives of $V$ are evaluated on $R_{0}(t)$. Varying with respect to $\delta R$ gives
The angular fluctuation is not canonically normalized. Defining the canonical axial fluctuation
$\delta\psi(t,\bm{x})\equiv f(R_{0}(t))\,\delta\theta(t,\bm{x}),$ (62)
the $\delta\theta$ equation can be rewritten in Klein–Gordon form,
with time-dependent effective mass
$m_{\psi}^{2}(t)=\frac{f^{\prime}(R_{0})}{f(R_{0})}V_{,R}(R_{0})-\frac{f^{\prime\prime}(R_{0})}{f(R_{0})}\,\dot{R}_{0}^{2}.$ (64)
Thus, in the vanishing-charge background, the internal geometry affects the linear dynamics through the axial perturbations, while the radial perturbation $\delta R$ is governed by the usual potential curvature $V_{,RR}(R_{0})$.

### v2

@@ Abstract
CMB limits on cold-dark-matter isocurvature are often interpreted as excluding the simultaneous realization of high-scale inflation and large QCD axion decay constants in pre-inflationary Peccei–Quinn (PQ) scenarios. We show that this conclusion can be evaded by exploiting field-space geometry. For a minimal complex PQ scalar with a $U(1)$-symmetric potential and a nonlinear sigma-model kinetic term $d\sigma^{2}=dR^{2}+f^{2}(R)\,d\theta^{2}$, the observable axion fluctuation is $\delta\theta\sim H_{\rm inf}/f(R)$, so an enhanced effective decay constant $f(R)$ suppresses isocurvature without explicit PQ breaking, extreme radial displacements, or additional couplings. Specializing to a hyperbolic metric $f(R)\propto\sinh(R/L)$ with curvature scale $L$, we find that for $R\gtrsim L$ the exponential geometric-factor $f(R)$ strongly suppresses CMB-scale isocurvature while producing a characteristic blue-tilted spectrum. The same geometry induces a time-dependent $\mathcal{O}(H_{\rm inf})$ effective mass for the canonical axial mode even during radial slow-roll, which controls the scale-dependence of isocurvature. As a result, inflationary Hubble scales as large as $H_{\rm inf}\sim 10^{13}\,\mathrm{GeV}$ can be compatible with $f_{a}\sim 10^{14}$$10^{16}\,\mathrm{GeV}$, reopening parameter space usually regarded as excluded. We present numerical benchmarks and a semi-analytic template that relates the scale-dependence of isocurvature to the geometric lever arm $R/L$, providing a direct phenomenological probe on PQ field-space geometry.
@@ I Introduction
In this work we present a minimal, $U(1)$ conserving alternative based on field-space geometry. Curved internal field-space metrics are well known to generate effective masses and characteristic scale-dependence for entropic/isocurvature modes in covariant sigma-model systems [8, 46, 14, 35, 24]. In this Letter, we apply this mechanism to a minimal PQ field (in particular, a quartic PQ potential with spontaneously-broken $U(1)_{\rm PQ}$), and show that a hyperbolic field-space metric exponentially suppresses the observable angular fluctuation without explicit PQ breaking. This dynamically suppresses CMB-scale axion isocurvature, while the evolving mass produces a distinctive blue-tilted isocurvature spectrum on smaller scales. As a result, viable parameter space with $H_{\rm inf}\sim 10^{13}\,{\rm GeV}$ and large $f_{a}$ (e.g. $\lesssim 10^{16}\,{\rm GeV}$) can be reopened without requiring the extreme displacements $R_{\rm inf}/f_{a}\sim 10^{4}$$10^{5}$ or tiny quartic coefficients $\lesssim\mathcal{O}(10^{-16})$. We further provide a compact phenomenological template relating the isocurvature spectral index to the field-space curvature scale, thereby promoting measurements of isocurvature scale-dependence to a direct probe of PQ field-space geometry and enabling targeted searches in future analyses.
The paper is organized as follows. Section II reviews the axion isocurvature constraint, Sec. III introduces the hyperbolic PQ field-space framework, and Sec. IV presents benchmarks and resulting spectra. We conclude in Sec. V. Appendices A and B give a UV-motivated origin for the hyperbolic metric, and a semi-analytic data-facing template respectively.
@@ II The standard QCD axion isocurvature problem
$T_{\rm inst.RH}$ is the instantaneous reheating temperature and the orange solid line marks the upper boundary of the region for PQ restoration post-inflation. In flat-metric PQ models, realizing simultaneously $H_{\rm inf}\simeq 10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{16}\,{\rm GeV}$ typically forces $\kappa\gtrsim 10^{4}$, often into a regime where the spectator/EFT description is difficult to control. This motivates alternative suppression mechanisms.
@@ III Hyperbolic PQ field-space and geometric suppression
See Supplemental Material at [52] for derivations in this section. Details and derivations in this section are deferred to the Supplemental Material. We work in the vanishing-charge background $\dot{\theta}_{0}=0$, for which the homogeneous evolution is purely radial,
@@
Defining the canonically normalized angular fluctuation $\delta\psi\equiv f(R_{0})\,\delta\theta$, one finds
@@
The leading geometric contribution is enhanced by the curvature lever arm $\xi=R_{0}/L$, while the subleading negative term encodes the reduction of transverse masses from negative curvature when $\dot{R}_{0}\neq 0$. Unlike for a flat-metric, the axial mode can become $\mathcal{O}(H)$-heavy even during radial slow-roll. The time dependence generically produces blue-tilted isocurvature: modes exiting earlier (larger scales) are most suppressed, while later-exiting modes (smaller scales) are less suppressed. From Eq. (16), we also note that large radial velocities (underdamped motion) tend to destabilize the angular mode.
@@ IV Benchmark models and isocurvature spectrum
$\Delta_{\theta}^{2}(k)\equiv\frac{k^{3}}{2\pi^{2}}\,|\delta\theta_{k}|^{2},\qquad\delta\theta_{k}=\left.\frac{\delta\psi_{k}}{f(R_{0})}\right|_{\rm late-time}.$ (20)
@@
We consider three benchmark points B1-B3 (Tab. 1), chosen such that the initial curvature lever arm $\xi_{i}\equiv R_{i}/L$ is $\sim\mathcal{O}(10)$ and CMB-scale isocurvature is evaluated at a high $H_{\rm inf}\sim 10^{13}~\rm{GeV}$. The near-degeneracy of B1 and B3 is intentional: in $H_{\rm inf}$ units the spectator PQ system has an approximate rescaling symmetry,
@@ IV.1 Results
Fig. 2 shows the mass history for B1. The radial field slow-rolls toward its vacuum, while the hyperbolic metric induces a geometric axial mass. Initially $m_{\psi}^{2}\sim\mathcal{O}(H_{\rm inf}^{2})$ while $m^{2}_{R}\ll H_{\rm inf}^{2}$, so modes exiting early are damped. As $R_{0}$ decreases and $\dot{R}_{0}$ redshifts, $m_{\psi}^{2}/H_{\rm inf}^{2}$ falls and the damping weakens for later-exiting modes. In the regime $\delta=\xi(m_{R}/H_{\rm inf})^{2}\lesssim 1$, the dominant behavior is $m_{\psi}^{2}\simeq\xi\,m_{R}^{2}$, i.e. the angular mode is enhanced relative to the flat case by the curvature lever arm $\xi=R/L$, while for $\xi m_{R}^{2}\gtrsim 1$ the negative curvature term $\propto-\dot{R}^{2}$ reduces the effective axial mass.
@@
Fig. 3 shows the resulting conserved superhorizon $\Delta_{\mathcal{S}}^{2}(k)$. We take $k/(a_{i}H_{\rm inf})=10$ to correspond to the CMB pivot scale $k_{\ast}\simeq 0.002,{\rm Mpc}^{-1}$. Power is strongly suppressed on the largest CMB scales because those modes exit while $f(R)$ is exponentially enhanced, so that the observable fluctuation $\delta\theta=\delta\psi/f(R)$ is strongly suppressed at horizon exit. As $R$ slow-rolls during the subsequent radial evolution, $f(R(t_{k}))$ decreases and smaller scales are progressively less suppressed. The outcome is a blue-tilted isocurvature spectrum with running over the range of modes that probe the evolution of $R$ (and hence $f$), approaching a quasi-plateau once the angular mode becomes effectively light. Equivalently, this behavior can be described in terms of the time-dependent effective mass $m_{\psi}^{2}(t)$ during radial slow-roll, which governs the superhorizon evolution of $\delta\psi$ while leaving the ratio $\delta\psi/f(R)=\delta\theta$ nearly conserved in the slow-roll regime outside the horizon.
@@
A key result is that hyperbolic geometry can reopen parameter space that is excluded in the flat-metric case. In our benchmarks, we focus on $H_{\rm inf}=10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{14}\text{--}10^{16}\,{\rm GeV}$, as it lies near current upper limits and is the most demanding regime for isocurvature. In contrast, for $f(R)=R$ and comparable spectator trajectories, satisfying the CMB bound at $H_{\rm inf}=10^{13}\,{\rm GeV}$ typically requires extreme radial excursion ($R_{i}/f_{a}\sim 10^{4}\text{--}10^{5}$), often implying super-Planckian displacements.
@@
In the slow-roll geometric domain, the hyperbolic metric ties the axial mass to the radial background through $\xi\equiv R/L$. For quartic PQ slow-roll, we give the following scale-dependent mass template (see App. B)
@@
where $k_{m}$ is a matching scale at the onset of the geometric regime and $\xi_{m}\equiv R_{m}/L$ controls the shape. For the benchmark scenarios, a convenient prescription is to fix $k_{m}$ by $n_{\rm iso}(k_{m})=2$, equivalently $m_{\psi}^{2}(k_{m})/H_{\rm inf}^{2}=5/4$, so that for $k\gtrsim k_{m}$ the entire running is governed by the single curvature parameter $\xi_{m}$, which labels the family of trajectories.
The salient point is the functional form. In the hyperbolic regime one has $m_{\psi}^{2}\propto\xi\,m_{R}^{2}\propto R\,m_{R}^{2}$, and for a quartic slow-roll trajectory this gives
$m_{\psi}^{2}(N)\propto(N-N_{m})^{-3/2}\,\leftrightarrow\,m_{\psi}^{2}(k)\propto\left[\ln\!\left(\frac{k}{k_{m}}\right)\right]^{-3/2}.$ (25)
@@ V Conclusion
We have shown that the intrinsic geometry of the PQ field manifold can provide a minimal and symmetry-preserving solution to the standard axion isocurvature problem. For a hyperbolic PQ target-space metric, the angular fluctuation acquires a geometric, time-dependent effective mass during inflation, even in the absence of explicit PQ breaking in the axion potential. The same geometry exponentially enhances $f(R)$, thus suppressing CMB-scale isocurvature while yielding a new mechanism to generate blue-tilted spectrum even during radial slow-roll. In benchmark realizations with $H_{\rm inf}=10^{13}\,{\rm GeV}$, the mechanism reopens parameter space at large $f_{a}$ that is typically excluded in flat field-space.
Beyond enlarging the viable QCD-axion window for high-scale inflation, our results highlight a broader point: measurements of isocurvature scale-dependence can directly probe field-space curvature. We provided a compact phenomenological template that ties the running of the isocurvature tilt to a single curvature-controlled parameter, enabling model-agnostic tests of the mechanism and straightforward incorporation into future analyses. Several broad directions are motivated by this perspective. On the theory side, it will be important to embed controlled UV realizations of hyperbolic metrics to the PQ model, and extend the framework to more general target spaces (non-constant curvature and/or non-diagonal metrics). On the dynamics side, nontrivial background motion will be particularly interesting (such as $\dot{\theta}\neq 0$ or post-inflationary evolution). Hyperbolic geometry can naturally support large angular velocities, making contact with rotating-field mechanisms such as Affleck-Dine dynamics in baryogenesis ([2]), axiogenesis [22] and kinetic-misalignment in [21], and more generally with additional signatures beyond the power spectrum [51, 28]. Finally, the generic enhancement of power on sub-CMB scales motivates systematic confrontation with small-scale probes, including Lyman-$\alpha$ forest [9], UV luminosity functions of high-$z$ galaxies [57], 21-cm measurements [45], and strong-lensing substructure [23, 56], thereby enabling targeted searches for scale-dependent isocurvature as a window onto PQ field-space geometry.
@@ Appendix A An example of hyperbolic PQ field-space
In our setup the PQ sector is described by a single complex scalar $\Phi$. At the level of a low-energy effective theory, the most general two-derivative kinetic term takes the form of a nonlinear sigma model,
@@
where $\phi^{a}$ are the two real components of $\Phi$. The PQ symmetry is realized as a $U(1)$ isometry of the target space (shifts of the angular coordinate). We now give a concrete example showing how a hyperbolic geometry can arise.